OpenStudy (anonymous):

Please help me! The graph of f(x)= 1/(x^2-c) has a vertical asymptope at x=3. Find C.

OpenStudy (asnaseer):

If a function has a vertical asymptote at a particular value of x, it means that its value if infinite at that point. so, how can this equation: 1/(x^2-c) become infinite in value - can you guess what the condition must be?

OpenStudy (asnaseer):

*value is infinite

OpenStudy (anonymous):

any number greater than 0?

OpenStudy (asnaseer):

if 1/z is infinite, then what does that tell you about z?

OpenStudy (asnaseer):

i.e. what number can you divide by in order to infinity?

OpenStudy (anonymous):

1?

OpenStudy (asnaseer):

1/1 = infinity?

OpenStudy (anonymous):

I have no clue...

OpenStudy (asnaseer):

ok, think about this - what is 1/0.5 = ?

OpenStudy (anonymous):

2

OpenStudy (asnaseer):

good. now what is 1/0.05 = ?

OpenStudy (anonymous):

20

OpenStudy (asnaseer):

so - can you see that as the denominator becomes smaller then 1 divided by that number becomes bigger

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so think about this - what is the smallest value that the denominator could become?

OpenStudy (asnaseer):

0.5, 0.05, 0.005, 0.0005, ..., what value will we tend towards?

OpenStudy (anonymous):

.0005

OpenStudy (asnaseer):

if we keep smaking the denominator smaller and smaller - what will it eventually become?

OpenStudy (asnaseer):

*making

OpenStudy (anonymous):

infinate

OpenStudy (asnaseer):

no - that is making the denominator bigger and bigger

OpenStudy (anonymous):

all I am getting from this is that the answer needs to be .0000000000000000000000000000000000005

OpenStudy (asnaseer):

what is the smallest positive number you can think of?

OpenStudy (anonymous):

1

OpenStudy (asnaseer):

including zero

OpenStudy (anonymous):

0

OpenStudy (asnaseer):

yes, so 1/0 will be infinity

OpenStudy (asnaseer):

so, in your equation, you need:\[x^2-c=0\]at x=3

OpenStudy (anonymous):

so C=0?

OpenStudy (asnaseer):

first replace x^2 by 3^2 - what do you get?

OpenStudy (anonymous):

why are we replacing that and where did you get that option

OpenStudy (asnaseer):

your question says that there is an asymptote at x=3

OpenStudy (asnaseer):

so we need to replace x with 3 in the equation for f(x)

OpenStudy (anonymous):

so is c zero or not?

OpenStudy (asnaseer):

correct - c is not zero

OpenStudy (anonymous):

so what is c?

OpenStudy (asnaseer):

follow the steps I have outlined above and you should be able to solve for c.

OpenStudy (anonymous):

i appreciate you taking forever to explain this to me, but i have about 500 more problems to do and not much time to do them.

OpenStudy (anonymous):

@asnaseer i think that asymptope means approcing to a value but not reaches to it then it is not necessary to be infinity :)

OpenStudy (asnaseer):

@Aperogalics The value of f(x) at x=3 will be infinite if f(x) has a vertical asymptote at x=3

OpenStudy (anonymous):

no @asnaseer i m not going with this question i m talking about asymptote only :)

OpenStudy (anonymous):

general :)

OpenStudy (asnaseer):

of course in general asymptotes do not have to be vertical

OpenStudy (anonymous):

I just need to know C

OpenStudy (asnaseer):

@stevenchannell substitute x=3 into this equation:\[x^2-c=0\]and then solve to find c.

OpenStudy (anonymous):

no my mean to say not by vertical or horizontal or inclined but reaching to infinity is not the case which mean asymptote:)

OpenStudy (asnaseer):

oh I see what you are trying to say now - yes - you are right :)

OpenStudy (anonymous):

c=9?

OpenStudy (asnaseer):

perfect! @stevenchannell

OpenStudy (asnaseer):

my advise to you would be that it is better to take things slowly and understand the subject matter rather than trying to rush through lots of questions. :)

OpenStudy (anonymous):

yes I understand that, but all my work is online work and I will be leaving out of town for the next week